# SMILE: Zero-Shot Sparse Mixture of Low-Rank Experts Construction From Pre-Trained Foundation Models

### v1

@@ Abstract
Deep model training on extensive datasets is increasingly becoming cost-prohibitive, prompting the widespread adoption of deep model fusion techniques to leverage knowledge from pre-existing models. From simple weight averaging to more sophisticated methods like AdaMerging, model fusion effectively improves model performance and accelerates the development of new models. However, potential interference between parameters of individual models and the lack of interpretability in the fusion progress remain significant challenges. Existing methods often try to resolve the parameter interference issue by evaluating attributes of parameters, such as their magnitude or sign, or by parameter pruning. In this study, we begin by examining the fine-tuning of linear layers through the lens of subspace analysis and explicitly define parameter interference as an optimization problem to shed light on this subject. Subsequently, we introduce an innovative approach to model fusion called zero-shot Sparse MIxture of Low-rank Experts (SMILE) construction, which allows for the upscaling of source models into an MoE model without extra data or further training. Our approach relies on the observation that fine-tuning mostly keeps the important parts from the pre-training, but it uses less significant or unused areas to adapt to new tasks. Also, the issue of parameter interference, which is intrinsically intractable in the original parameter space, can be managed by expanding the dimensions. We conduct extensive experiments across diverse scenarios, such as image classification and text generalization tasks, using full fine-tuning and LoRA fine-tuning, and we apply our method to large language models (CLIP models, Flan-T5 models, and Mistral-7B models), highlighting the adaptability and scalability of SMILE. For full fine-tuned models, about 50% additional parameters can achieve around 98-99% of the performance of eight individual fine-tuned ViT models, while for LoRA fine-tuned Flan-T5 models, maintaining 99% performance with only 2% extra parameters. Code is available at https://github.com/tanganke/fusion_bench.
@@ 2 Rethinking Model Fine-Tuning From a Subspace Perspective
$\langle x_{ab},x_{cd}\rangle=\operatorname{tr}\left(u_{a}v_{b}^{T}(u_{c}v_{d}^{T})^{T}\right)=\operatorname{tr}\left(u_{a}v_{b}^{T}v_{d}u_{c}^{T}\right)\in\mathbb{R}.w$ (1)
@@ 5 Experiments
In this section, we evaluate the effectiveness of the proposed SMILE on a variety of setups, including image classification and text generalization tasks, as well as full fine-tuning and LoRA fine-tuning. Detailed information about the fine-tuned models is in Appendix C. We compare our method with several SOTA model fusion techniques, including Simple Averaging (Wolf et al., 2019b), Fisher merging (Matena and Raffel, 2022), RegMean (Jin et al., 2022), Task Arithmetic (Ilharco et al., 2022), Ties-Merging (Yadav et al., 2023), AdaMerging (Yang et al., 2024c), and WEMoE (Tang et al., 2024c). To further demonstrate the scalability of SMILE upscaling, we also conduct experiments using off-the-shelf large language models fine-tuned from Mistral-7B-v0.1. Our code implementation is based on FusionBench (Tang et al., 2024a).
@@ 5.1 Multi-Task Model Fusion on Open-Vocabulary Image Classification Tasks
5.2 Multi-Task Model Fusion on Text Generalization Tasks
We further evaluate SMILE on text generalization tasks using Flan-T5-base models  [^3] https://huggingface.co/google/flan-t5-base, which are fine-tuned on eight tasks from the GLUE benchmark (Wang et al., 2018). We use two different fine-tuning strategies: full fine-tuning and LoRA fine-tuning with $r_{LoRA}=16$. We present the results in Tables 4 and 5, for full fine-tuned models and LoRA fine-tuned models, respectively. For fully fine-tuned models, SMILE consistently outperforms other fusion methods across all eight tasks. With just 1.52 times the parameters of a single model, SMILE (2: $k_{gate}=8,k=32$) achieves 99.0% of the individual model performance with 1.52 times the parameters of a single model. In the LoRA fine-tuned scenario, SMILE maintains strong performance with minimal parameter increase (1.02 times). It achieves 99.3% of the individual model performance, significantly surpassing other multi-task model fusion methods.
@@ 6 Related Work
Mixture of Experts. The concept of Mixture of Experts (MoE) is first introduced by Jacobs et al. (1991), involving training multiple specialized models. This concept has gained significant attention in recent years (Jiang et al., 2024; Dai et al., 2024), with much of the innovation focusing on routing mechanisms and expert design. Much innovation revolves around the design of more efficient routers. For example, the Switch Transformer (Fedus et al., 2022b) selects only the top expert for each token, simplifying the process and improving scalability. Similarly, (Lewis et al., 2021) use a linear assignment to optimize token-expert affinities, ensuring an equal spread of tokens among experts. For detailed reviews on MoE, see (Fedus et al., 2022a) and for MoE in the context of model merging, refer to (Yadav et al., 2024).
@@ 7 Conclusion, Limitations, and Future Work
Understanding which subspaces contribute most to task-specific performance could lead to more targeted and efficient fine-tuning strategies, potentially focusing on updating specific parts of the model while leaving others intact. Additionally, this approach might be applied to other areas, like multi-modal large language models, where different types of data (modalities) are treated as separate experts. Furthermore, it would be worth exploring how SMILE can manage multi-objective optimization by adjusting the importance of different routing weights. What’s more, develop methods to dynamically adjust the number of experts $K$ per token based on the input, potentially improving efficiency without sacrificing performance.
@@ Appendix D Hyperparameter Analysis
(c) Heatmap of average performance.
(d) Line plot of average performance.

### v2

@@ Abstract
Deep model training on extensive datasets is increasingly becoming cost-prohibitive, prompting the widespread adoption of deep model fusion techniques to leverage knowledge from pre-existing models. From simple weight averaging to more sophisticated methods like AdaMerging, model fusion effectively improves model performance and accelerates the development of new models. However, potential interference between parameters of individual models and the lack of interpretability in the fusion progress remain significant challenges. Existing methods often try to resolve the parameter interference issue by evaluating attributes of parameters, such as their magnitude or sign, or by parameter pruning. In this study, we begin by examining the fine-tuning of linear layers through the lens of subspace analysis and explicitly define parameter interference as an optimization problem to shed light on this subject. Subsequently, we introduce an innovative approach to model fusion called zero-shot Sparse MIxture of Low-rank Experts (SMILE) construction, which allows for the upscaling of source models into an MoE model without extra data or further training. Our approach relies on the observation that fine-tuning mostly keeps the important parts from the pre-training, but it uses less significant or unused areas to adapt to new tasks. Also, the issue of parameter interference, which is intrinsically intractable in the original parameter space, can be managed by expanding the dimensions. We conduct extensive experiments across diverse scenarios, such as image classification and text generation tasks, using full fine-tuning and LoRA fine-tuning, and we apply our method to large language models (CLIP models, Flan-T5 models, and Mistral-7B models), highlighting the adaptability and scalability of SMILE. For full fine-tuned models, about 50% additional parameters can achieve around 98-99% of the performance of eight individual fine-tuned ViT models, while for LoRA fine-tuned Flan-T5 models, maintaining 99% performance with only 2% extra parameters. Code is available at https://github.com/tanganke/fusion_bench.
@@ 2 Rethinking Model Fine-Tuning From a Subspace Perspective
$\langle x_{ab},x_{cd}\rangle=\operatorname{tr}\left(u_{a}v_{b}^{T}(u_{c}v_{d}^{T})^{T}\right)=\operatorname{tr}\left(u_{a}v_{b}^{T}v_{d}u_{c}^{T}\right)\in\mathbb{R}.$ (1)
@@ 5 Experiments
In this section, we evaluate the effectiveness of the proposed SMILE on a variety of setups, including image classification and text generation tasks, as well as full fine-tuning and LoRA fine-tuning. Detailed information about the fine-tuned models is in Appendix C. We compare our method with several SOTA model fusion techniques, including Simple Averaging (Wolf et al., 2019b), Fisher merging (Matena and Raffel, 2022), RegMean (Jin et al., 2022), Task Arithmetic (Ilharco et al., 2022), Ties-Merging (Yadav et al., 2023), AdaMerging (Yang et al., 2024c), and WEMoE (Tang et al., 2024c). To further demonstrate the scalability of SMILE upscaling, we also conduct experiments using off-the-shelf large language models fine-tuned from Mistral-7B-v0.1. Our code implementation is based on FusionBench (Tang et al., 2024a).
@@ 5.1 Multi-Task Model Fusion on Open-Vocabulary Image Classification Tasks
5.2 Multi-Task Model Fusion on Text Generation Tasks
We further evaluate SMILE on text generation tasks using Flan-T5-base models  [^3] https://huggingface.co/google/flan-t5-base, which are fine-tuned on eight tasks from the GLUE benchmark (Wang et al., 2018). We use two different fine-tuning strategies: full fine-tuning and LoRA fine-tuning with $r_{LoRA}=16$. We present the results in Tables 4 and 5, for full fine-tuned models and LoRA fine-tuned models, respectively. For fully fine-tuned models, SMILE consistently outperforms other fusion methods across all eight tasks. With just 1.52 times the parameters of a single model, SMILE (2: $k_{gate}=8,k=32$) achieves 99.0% of the individual model performance with 1.52 times the parameters of a single model. In the LoRA fine-tuned scenario, SMILE maintains strong performance with minimal parameter increase (1.02 times). It achieves 99.3% of the individual model performance, significantly surpassing other multi-task model fusion methods.
@@ 5.3 Spare Mixture of Low-Rank Experts Analysis
(a) average performance
(b) normalized parameters count
(c) average performance
(d) normalized parameters count
@@ 6 Related Work
Mixture of Experts. The concept of Mixture of Experts (MoE) is first introduced by Jacobs et al. (1991), involving training multiple specialized models. This concept has gained significant attention in recent years (Jiang et al., 2024; Dai et al., 2024), with much of the innovation focusing on routing mechanisms and expert design. Much innovation revolves around the design of more efficient routers. For example, the Switch Transformer (Fedus et al., 2022b) selects only the top expert for each token, simplifying the process and improving scalability. Similarly, (Lewis et al., 2021) use a linear assignment to optimize token-expert affinities, ensuring an equal spread of tokens among experts. In (Ostapenko et al., 2024), the authors propose to build a library of LoRA adapters using model-based clustering and build a MoE to select the most relevant adapters based on input without retraining. For detailed reviews on MoE, see (Fedus et al., 2022a), and for MoE in the context of model merging (MoEerging), refer to (Yadav et al., 2024).
@@ 7 Conclusion, Limitations, and Future Work
Understanding which subspaces contribute most to task-specific performance could lead to more targeted and efficient fine-tuning strategies, potentially focusing on updating specific parts of the model while leaving others intact. Additionally, this approach might be applied to other areas, like multi-modal large language models, where different types of data (modalities) are treated as separate experts. Furthermore, it would be worth exploring how SMILE can manage multi-objective optimization by adjusting the importance of different routing weights. Moreover, develop methods to dynamically adjust the number of experts $K$ per token based on the input, potentially improving efficiency without sacrificing performance.
@@ Appendix D Hyperparameter Analysis
(c) Heatmap of normalized parameter count.
(d) Line plot of normalized parameter count.